OpenStudy (anonymous):

A cylindrical water tank has a diameter of 6 ft and a height of 8 ft. What is the exact volume of the tank? A. 24π ft³ B. 48π ft³ C. 72π ft³ D. 288π ft³

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

volume would be = pi r^2 l= pi 3^2 *8=72 pi

OpenStudy (whpalmer4):

\[V = \pi r^2 h\]\[r = \frac{d}2\]\[r = \frac{6}{2} = 3\]\[h = 8\]\[V = \pi (3)^2(8) = \]

OpenStudy (anonymous):

so the ans Is C

OpenStudy (whpalmer4):

Yes, C is correct. How did you arrive at B? It's valuable to understand your mistakes, so you can avoid repeating them in the future.